OpenStudy (anonymous):

the distance travelled by a body falling freely from in the 1st, 2nd and 3rd seconds, are in the ratio a)1:2:3 b)1:3:5 c)1:4:9 d)3:2:1

OpenStudy (shamim):

i think this is a problem of physics

OpenStudy (shamim):

anyway let the initial velocity\[u=0\]

OpenStudy (shamim):

the free falling body travel the distance in first 1 second is\[h _{1}=ut+\frac{ 1 }{ 2 }g t ^{2}=0 \times 1+\frac{ 1 }{ 2 }g \times 1^{2}=\frac{ 1 }{ 2 }g\]

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

i got the result \[1:3:5\]

OpenStudy (shamim):

i wanna show procedure if u want

OpenStudy (anonymous):

since the average of things falling is 9.8m/s2, then according to your question, the answer would be: 9.8 : 39.2 : 88.2 so. the answer would be: c) 1:4:9